At the end of the day, a storekeeper had $1638 in the cash register, including both the sale of goods and the sales tax of 4%. What amount of sales tax had been collected?
Answer by oberobic(2304)

t = sales
.04t = sales tax
1.04t = total money collected
.
1.04t = 1638
.
t = 1638/1.04
.
t = 1575 sales
.
.04t = 1638-1575 = 63
.
Always check your work.
.
.04(1575) = 63
Correct.
.
Re-read the question to be sure you answer it.
What amount of sales tax was collected? $63.
.
Done.